OpenStudy (anonymous):

Find SP, if MP = Rs. 70 and commission =10 percent

OpenStudy (mayankdevnani):

do you know the formula for finding sp????

OpenStudy (mayankdevnani):

@msingh

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

tell

OpenStudy (mayankdevnani):

sorry....plz write...same thing

OpenStudy (anonymous):

SP=CP-loss

OpenStudy (mayankdevnani):

is cp equal to mp???

OpenStudy (mayankdevnani):

@KingAditya this type of question is for you????

OpenStudy (mayankdevnani):

carry on!!! @msingh

OpenStudy (anonymous):

SP=70+loss

OpenStudy (anonymous):

I DONT LIKE COMMERCIAL MATHEMATICS...@mayankdevnani

OpenStudy (mayankdevnani):

why you add 70....to find the sp

OpenStudy (mayankdevnani):

you are correct!!! sir

OpenStudy (mayankdevnani):

to find the price of commision .......10 %* 70 to give the exactly price of commssion!!!! got it @msingh

OpenStudy (anonymous):

k

OpenStudy (mayankdevnani):

\[\frac{10}{100} \times 70=?????\]

OpenStudy (mayankdevnani):

solve it????

OpenStudy (anonymous):

7

OpenStudy (mayankdevnani):

correct!!! it means commission=Rs.7 !!! correct???

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

so put this in formula , we get, SP=MP+Rs.70 and MP=Rs.70 so, SP=Rs.70+Rs.7=??????? what we get????

OpenStudy (anonymous):

how you write SP=MP +70

OpenStudy (mayankdevnani):

we can write it because when we talk about commission ....we add to mp to get sp!!!!

OpenStudy (anonymous):

you didn't get what i m saying plse check your steps again

OpenStudy (mayankdevnani):

hey!!! this formula is different ..... SP=MP+COMMISSION... @got it

OpenStudy (anonymous):

wahi to what i m trying to say is that u have wrote SP=MP+70 inplace of SP=MP+7